# Match, Compare, or Select? An Investigation of Large Language Models for Entity Matching

### v2

@@ 3.5 Compound Entity Matching Framework
Our ComEM framework integrates the advantages of different strategies through a filtering then identifying pipeline. It first utilizes the local matching or comparing strategy to rank potential matches for preliminary screening, which can effectively mitigate the position bias and the long context requirement of the selecting strategy. It then utilizes the global selecting strategy to incorporate record interactions for fine-grained optimization, which can effectively mitigate the consistency ignorance of the matching strategy. Therefore, ComEM is able to strike a balance between entity matching requirements and current LLM capabilities, achieving significant performance improvements.
By integrating LLMs of different sizes, our ComEM framework can also effectively reduce the cost of LLM invocations for entity matching. In practice, direct use of commercial LLMs is expensive because entity matching is a computationally intensive task. ComEM delegates a significant part of the computation to medium-sized LLMs. Experimental results show that the ranking process can be performed well by using open-source medium-sized (3B~11B) LLMs (§ 4.4). As a result, the proper integration of LLMs not only improves the performance of entity matching but also reduces the cost for practical application.
Figure 2: Illustration of ComEM framework. It first filters candidate records by matching or comparing strategies and then identifies the match via selecting strategy.
4 Experiments
@@
Table 2: Statistics of experimental datasets.
Method Metric AB AG DA DS IM IV TT WA Mean Cost
P 71.43 38.00 86.71 71.19 84.44 57.81 67.97 73.49 68.88
R 50.00 75.00 95.67 84.00 85.00 95.00 87.00 40.67 76.54
Sudowoodo F1 58.82 50.45 90.97 77.06 84.72 71.88 76.32 52.36 70.32 1.11
P 40.41 35.54 65.78 64.63 95.08 68.75 65.28 35.62 58.89
R 91.33 59.00 98.67 81.00 58.00 55.00 94.00 88.33 78.17
Matching F1 56.03 44.36 78.93 71.89 72.05 61.11 77.05 50.77 64.02 4.52
P 81.69 65.31 85.60 82.74 96.55 84.82 88.93 71.26 82.11
R 77.33 42.67 69.33 54.33 46.67 31.67 85.67 60.33 58.50
Comparing F1 79.45 51.61 76.61 65.59 62.92 46.12 87.27 65.34 66.86 11.75
P 74.08 58.13 81.34 73.89 89.41 84.07 77.18 72.95 76.38
R 87.67 70.33 97.33 88.67 95.67 82.67 91.33 89.00 87.83
@@ 4.1 Experimental Setup
Datasets.  We focused on the common record linkage that has many open-access datasets. Specifically, we used 8 clean-clean ER datasets collected by pyJedAI Nikoletos et al. (2022). Table 2 shows the statistics of these datasets. We adapted them to fit the problem formulation and to support our experiments. For each dataset with two record collections D1 and D2, we applied the SOTA blocking method Sparkly Paulsen et al. (2023) to retrieve 10 potential matches from D2 for each record in D1. The recall@10 of Sparkly on all datasets ranges from 86.57% to 99.96%, demonstrating its effectiveness in retrieving potential matches. In this way, we are able to investigate and evaluate different strategies under the real ER pipeline.
Baseline.  Except for the pairwise matching strategy, we also compare the STOA self-supervised learning method, Sudowoodo Wang et al. (2023), which reduces the need for supervision through contrastive learning and pseudo-labeling.
Evaluation Metrics.  Consistent with prior studies, we report F1, Precision, and Recall on record pairs. We also report the cost (\$) of LLM invocations [^2] The inference or training cost is estimated based on the hourly price of the cloud NVIDIA A40..
Implementation Details. [^3] We have provided the full code, including blocking and sampling in the Supplementary Material for reproducibility. We used ChatGPT (gpt-3.5-turbo-0613) as the main LLM for strategy analysis. We also examined the effect of 8 open-source instruction-tuned LLMs, including Llama-3-8B (AI@Meta, 2024), Qwen2-7B (Bai et al., 2023), Mistral-7B (Jiang et al., 2023), Mixtral-8x7B (Jiang et al., 2024), Flan-T5-XXL (Chung et al., 2022), Flan-UL2 (Tay et al., 2023) and Solar-10.7B (Kim et al., 2023). The specific prompts can be found in Appendix A, with the generation temperature of all LLMs set to 0 for reproducibility. For each dataset, we sampled 400 records from record collection D1 for evaluation, 300 of which had matches, and formed 4000 pairs of records with their potential matches obtained by Sparkly from record collection D2. The remaining record pairs (unsampled records and their potential matches) are used for model training or in-context learning. For Sudowoodo, we used its official implementation [^4] https://github.com/megagonlabs/sudowoodo to train models on 500 pairs. For in-context learning, we select 100 record pairs and follow Peeters and Bizer (2023a) to retrieve 3 positives and 3 negatives as few-shot examples. Since the comparing strategy produces only relative orders, we applied the matching strategy to the top 1 candidate after bubble sort ranking. In ComEM, we used Flan-T5-XL to rank candidates with the matching strategy and kept the top 4 candidates for selection.
@@ 4.2 Main Results
Finding 1. Incorporating record interactions is essential for LLM-based entity matching. As shown in Table 3, entity matching performance increases as the level of record interaction increases. The comparing strategy outperforms the independent matching strategy by an average of 2.84% F1 score, and the selecting strategy further improves the F1 score by up to 14.74% over the comparing strategy. The optimal selecting strategy is 11.28% higher in F1 than the SOTA self-supervised learning method. The advantages of the comparing and selecting strategies over the matching strategy are also evident across different LLMs in Figure 5. To further verify that these improvements are due to the strategy, we perform 6-shot matching, ensuring that the number of records is consistent with the selecting strategy. As shown in Figure 3, the selecting strategy still has a significant F1 advantage over 6-shot matching. These results highlight the effectiveness of our proposed strategies and open new avenues for LLM-based entity matching.
Finding 2. By integrating the advantages of different strategies and LLMs, ComEM can accomplish EM more effectively and cost-efficiently. As shown in Table 3 and Figure 3, compared to the optimal selecting strategy using ChatGPT, ComEM achieves 4% F1 improvement while spending less. The filtering and identifying pipeline improves precision considerably (6.7%) without sacrificing high recall of the selecting strategy. These results reveal that integrating multiple strategies can complement single strategies and mitigate the position bias of the selecting strategy in long contexts. However, using a single powerful but costly commercial LLM to complete the entire pipeline obscures the cost efficiency of the selecting strategy. By introducing a medium-sized LLM for preliminary filtering, ComEM improves performance while spending less than direct selection. As a result, ComEM underscores the importance of task decomposition and LLM composition, illuminating an effective route for compound entity matching using LLMs.
Figure 4: F1 score w.r.t. matched record positions.
@@ 4.3 Analysis of Strategies
Finding 3. The selecting is the most cost-effective strategy for LLM-based entity matching. Monetary cost is also an important factor when interfacing LLMs for EM in practice, as it is computationally intensive. As shown in Table 3, the selecting strategy costs less than half of the matching strategy. This is because the selecting strategy saves $n-1$ times of repeatedly inputting anchor records and task instructions into LLMs. The comparing strategy, however, considers two potential matches at a time and interfaces the LLM twice, making its cost more than twice that of the matching strategy. Therefore, the selecting strategy stands out for its effectiveness and efficiency.
Finding 4. Strategies that incorporate multiple records suffer from the position bias of LLMs. As shown in Figure 4, the performance of the comparing and selecting strategies decreases significantly as the position of the matching records moves down in the candidate list. For the comparing strategy optimized with bubble sort, matching records cannot be ranked at the top if there is any incorrect comparison. The selecting strategy also drops about 10% in F1, probably due to the limited long context understanding of the LLM. Therefore, the position bias of LLMs restricts the performance and generality of the comparing and selecting strategies.
Figure 5: Effect of open-source LLMs on different strategies.
@@ 4.4 Effect of LLMs
Finding 5. There is no single LLM that is uniformly dominant across all strategies. Figure 5 shows the efficacy of proposed strategies for open-source LLMs, with detailed results in Appendix B. We can see that the F1 scores of the matching, comparing, and selecting strategies for different LLMs mostly fall between 50%~70%, 60%~80%, and 70%~80%, respectively. In general, similar to ChatGPT, the comparing strategy is better than the matching strategy, while the selecting strategy is further better than the comparing strategy. The consistent performance between strategies confirms the effectiveness of incorporating record interactions in these ways. Concretely, some chat LLMs such as Llama3-8B and Mistral-7B produce numerous false positives and thus perform poorly with the matching strategy. Nevertheless, they achieve significant improvement and satisfactory performance by comparing and selecting strategies, respectively. Moreover, although Flan-T5-XXL and Flan-UL2 lag behind ChatGPT by about 4% F1 in the selecting strategy, we find that they perform quite well in the matching and comparing strategies. These task-tuned LLMs follow instructions better than chat-tuned LLMs, and can output only the requested labels instead of long-form responses, making it convenient to utilize label generation probabilities. In conclusion, there is a noticeable variance in the capabilities of different LLMs for a single strategy, and the efficacy of different strategies for a single LLM can also be significantly distinct.
@@
Finding 6. Matching strategy is better for ranking and filtering than comparing strategy. The superiority of Flan-T5 in the matching and comparing strategies leads us to explore the possibility of using it to rank and filter potential matches for the selecting strategy. As shown in Figure 6, the matching strategy outperforms the comparing strategy under different model parameter sizes, even though the latter performs $\mathcal{O}(n^{2})$ comparisons. The difference is small on Flan-T5-XL (3B) and Flan-T5-XXL (11B), but significant on smaller models. This may be due to the fact that these models are trained on many pairwise tasks, such as natural language inference and question answering, but few triple-wise tasks. Therefore, in terms of effectiveness and efficiency, the matching strategy is more suitable for ranking and filtering potential matches.
@@ 4.5 Ablation Study
We perform an ablation study on the number of candidate records for further identification. As shown in Figure 7, recall increases and precision decreases as the number of retained potential matches increases. Consistent with Figure 4, four is the sweet spot for the selecting strategy with current ChatGPT, which balances precision and recall well.
Figure 7: Average F1, precision, and recall w.r.t. number of candidate retained ($k$) for further selection. See Appendix C for detailed results on each dataset.
@@ 5 Conclusion
In this paper, we investigate three strategies for entity matching using LLMs to bridge the gap between local matching and global consistency of ER. Our research shows that incorporating record interactions is essential for LLM-based entity matching. By examining the effect of broad LLMs on these strategies, we further design a ComEM framework that integrates the advantages of multiple strategies and LLMs. The effectiveness and cost efficiency of ComEM highlight the importance of task decomposition and LLM composition, opening up new avenues for entity matching using LLMs.
@@
This study aims to investigate different strategies for LLM-based entity matching. We conducted thorough experiments with 1 commercial LLM and 8 open-source LLMs to provide a broad base for our analysis. The selection of models is based on considerations of popularity, availability, and cost. Future research could explore whether similar findings hold as LLMs evolve and how performance changes relative to our results.
Since LLMs have been trained on massive web data, they are likely to have seen similar and same records, or even some matching results, even though the labels of the matches are stored separately. Nevertheless, the performance of these strategies is relatively consistent across 9 LLMs and varies greatly for the same LLM when using different strategies, highlighting that data exposure is not the determining factor in their effectiveness. In the future, it will be valuable to evaluate LLM-based entity matching on new or non-public data.
The investigation of different strategies was conducted using basic zero/few-shot prompting, a simple and effective paradigm for applying LLMs. We could not ignore the role of potential advanced prompt engineering methods in improving the accuracy and robustness of LLMs. In addition, fine-tuning LLMs for better execution of different strategies is also a worthwhile direction.
@@
Appendix A Prompts
The prompts for various strategies of LLM-based entity matching used in this paper are presented in Table 4. To ensure fairness, the same prompts were used for all experimental LLMs. These prompts were constructed through a manual process of prompt engineering, which involved the testing and comparing of multiple variations to determine the most effective ones. In addition to the task description, we included specific response instructions such as “Answer "Yes" if they do and "No" if they do not” to guide the responses of LLMs. For in-context learning, prompts and labels were repeatedly inputted for each example, followed by the records to be matched. We post-processed the LLM responses to obtain the final predicted labels.
Matching
Do the two entity records refer to the same real-world entity? Answer "Yes" if they do and "No" if they do not.
Record 1: {{ record_left }}
Record 2: {{ record_right }}
Comparing
Which of the following two records is more likely to refer to the same real-world entity as the given record? Answer with the corresponding record identifier "Record A" or "Record B".
Given entity record: {{ anchor }}
Record A: {{ candidate_left }}
Record B: {{ candidate_right }}
Selecting
Select a record from the following candidates that refers to the same real-world entity as the given record. Answer with the corresponding record number surrounded by "[]" or "[0]" if there is none.
Given entity record: {{ anchor }}
Candidate records:{for candidate in candidates %}
[{{ loop.index }}] {{ candidate }}{endfor %}
Table 4: Specific prompts of different strategies. We use JinJa template syntax to display the placeholders for the anchor record and potential matches (candidates).
Appendix B Detailed Results of Open-Source LLMs under Different Strategies
We provide the detailed F1 scores of open-source LLMs under different strategies in Table 5. Among the 8 LLMs evaluated in our experiment, 6 achieve the best performance through the selecting strategy, and 2 achieve better performance through the comparing strategy. In summary, our proposed strategies are universally applicable across different LLMs for entity matching. We have observed that it is difficult to limit the output of many chat-tuned LLMs simply by prompts, which may affect their actual performance in entity matching. Therefore, how to calibrate the label probabilities from the long-form responses of LLMs is also important for performance improvement.

### v3

@@ 3.5 Compound Entity Matching Framework
Our ComEM integrates the advantages of different strategies through a filtering then identifying pipeline. It first utilizes the local matching or comparing strategy to rank potential matches for preliminary screening, which can effectively mitigate the position bias and the long context requirement of the selecting strategy. It then utilizes the global selecting strategy to incorporate record interactions for fine-grained optimization, which can effectively mitigate the consistency ignorance of the matching strategy. Therefore, ComEM is able to strike a balance between entity matching requirements and current LLM capabilities, achieving significant performance improvements.
By integrating LLMs of different sizes, ComEM can also effectively reduce the cost of LLM invocations for entity matching. In practice, direct use of commercial LLMs is expensive because entity matching is a computationally intensive task. ComEM delegates a significant part of the computation to medium-sized LLMs. Experimental results show that the ranking process can be performed well by using open-source medium-sized (3B~11B) LLMs (§ 4.4). As a result, the proper integration of LLMs not only improves the performance of entity matching, but also reduces the cost for practical application.
@@
Table 3: Statistics of experimental datasets. # denotes “number of”, D1 and D2 represent records from the 1st and 2nd sources, respectively. Attr and Pos refer to attributes of structured records and positive (matching) record pairs, respectively.
4 Experiments
AB AG DA DS IM IV TT WA Mean Cost
Supervised
Ditto Li et al. (2020) 77.34 63.79 93.80 90.02 97.06 78.59 87.15 57.75 80.69 0.29
HierGAT Yao et al. (2022) 75.51 64.45 98.01 89.47 97.69 77.73 85.34 78.55 83.34 1.10
Un/Self-supervised
ZeroER Wu et al. (2020) 32.66 22.14 99.32 84.14 43.32 0.50 53.76 61.52 49.67 /
Sudowoodo Wang et al. (2023) 58.82 50.45 90.97 77.06 84.72 71.88 76.32 52.36 70.32 0.63
GPT-3.5 Turbo
Matching Peeters et al. (2025) 56.03 44.36 78.93 71.89 72.05 61.11 77.05 50.77 64.02 4.52
Matching (6-shot) Peeters et al. (2025) 77.59 60.21 73.13 52.88 84.05 71.45 71.21 69.37 69.99 32.75
Comparing 79.45 51.61 76.61 65.59 62.92 46.12 87.27 65.34 66.86 11.75
Selecting 80.31 63.65 88.62 80.61 92.43 83.36 83.66 80.18 81.60 1.71
@@ 4.1 Experimental Setup
Datasets.  We focused on record linkage, a common form of entity resolution that identifies matching records between two data sources. Specifically, we used eight clean-clean ER datasets collected by pyJedAI Nikoletos et al. (2022). Table 3 shows the statistics of these datasets, where record collections D1 and D2 represent records from the first and second sources, respectively. For each dataset, we applied the SOTA blocking method Sparkly Paulsen et al. (2023) as preprocessing to retrieve 10 potential matches from D2 for each record in D1. The recall@10 of Sparkly on all datasets ranges from 86.57% to 99.96%, demonstrating its effectiveness in retrieving potential matches. We sampled 400 records from D1 for evaluation, 300 of which had matches, and formed 4,000 pairs of records by combining them with their potential matches from D2. To build training sets for learning-based methods, we further sampled 5,000 record pairs from the remaining records and their potential matches. Through this process, we constructed entity matching datasets that satisfied our formulation, with all methods evaluating on the same datasets after blocking to ensure a fair comparison.
Baseline.  We considered several SOTA methods as our baselines, including the unsupervised ZeroER Wu et al. (2020), the self-supervised Sudowoodo Wang et al. (2023), and the LLM-based matching strategy Peeters et al. (2025). For a comprehensive comparison, we also included two representative supervised methods, Ditto Li et al. (2020) and HierGAT Yao et al. (2022). [^3] We followed their open-source implementations and default parameters for reproduction.
Evaluation Metrics.  Consistent with prior studies, we report the F1 score as performance measure. We also report the cost (\$) of LLM invocations. For the compute cost of open-source LMs, we estimated it based on the training or inference time required and the hourly price of the cloud NVIDIA A40.
Implementation Details.  We used GPT-4o Mini (0718) and GPT-3.5 Turbo (0613) as the main LLMs for analysis. We also examined the effect of eight open-source instruction-tuned LLMs, including Llama-3.1-8B (Dubey et al., 2024), Qwen2-7B (Yang et al., 2024), Mistral-7B (Jiang et al., 2023), Mixtral-8x7B (Jiang et al., 2024), Command-R-35B, Flan-T5-XXL (Chung et al., 2024), Flan-UL2 (Tay et al., 2023) and Solar-10.7B (Kim et al., 2024). The specific prompts can be found in Appendix B, with the generation temperature of all LLMs set to 0 for reproducibility. For in-context learning, we retrieve 3 positives and 3 negatives as few-shot examples based on record similarity as Peeters et al. (2025). Since the comparing strategy produces only relative orders, we applied the matching strategy to the top 1 candidate after bubble sort comparing. In ComEM, we used Flan-T5-XL to rank candidates with the matching strategy and kept the top 4 candidates for selection.
@@ 4.2 Main Results
Finding 1. Incorporating record interactions is essential for LLM-based entity matching. As shown in Table 4, the performance of LLM-based entity matching increases with incorporating record interaction. The comparing strategy outperforms the independent matching strategy by an average of 10.7% F1 score, and the selecting strategy further improves the F1 score by an average of 5.32% over the comparing strategy. The advantages of the comparing and selecting strategies over the matching strategy are also evident across different LLMs in Figure 4. To further verify that these improvements are due to the strategy, we perform 6-shot matching, ensuring that the number of records is consistent with the selecting strategy. We can see that the selecting strategy still outperforms 6-shot matching by 7.97% in F1. Moreover, the proposed strategies enable LLM-based entity matching to surpass SOTA un/self-supervised methods and to be comparable to supervised methods that require extensive labeling data. These results highlight the effectiveness of our proposed strategies and open new avenues for LLM-based entity matching.
Finding 2. By integrating the advantages of different strategies and LLMs, ComEM can accomplish entity matching more effectively and cost-efficiently. As shown in Table 4, compared to the single comparing and selecting strategies, ComEM achieves 2~18% F1 improvements while spending less. The filtering and identifying pipeline improves precision considerably without sacrificing the high recall of the selecting strategy. These results reveal that integrating multiple strategies can complement single strategies and mitigate the position bias of the selecting strategy in long contexts. However, using a single powerful but costly commercial LLM to complete the entire pipeline obscures the cost efficiency of the selecting strategy. By introducing a medium-sized LLM for preliminary filtering, ComEM improves performance while spending less than direct selection. As a result, ComEM underscores the importance of task decomposition and LLM composition, illuminating an effective route for compound EM using LLMs.
Figure 3: F1 score w.r.t. matching record positions.
@@ 4.3 Analysis of Strategies
Figure 4: Effect of open-source LLMs on different strategies and ComEM.
@@
Finding 3. The selecting strategy is the most cost-effective strategy for LLM-based entity matching. Monetary cost is also an important factor when interfacing LLMs for entity matching in practice, as it is computationally intensive. As shown in Table 4, the selecting strategy costs less than half of the matching strategy. This is because the selecting strategy saves $n-1$ times of repeatedly inputting anchor records and task instructions into LLMs. The comparing strategy, however, considers two potential matches at a time and interfaces the LLM twice, making its cost more than twice that of the matching strategy. Therefore, the selecting strategy stands out for its effectiveness and efficiency.
Finding 4. Strategies that incorporate multiple records suffer from the position bias of LLMs. As shown in Figure 3, the performance of the comparing and selecting strategies varies significantly as the position of the matching records moves down in the candidate list. For the comparing strategy optimized with bubble sort, matching records cannot be ranked at the top if there is any incorrect comparison. The selecting strategy is also highly sensitive to the matching record positions, while ComEM can alleviate this. Therefore, the position bias of LLMs limits the performance and generality of the comparing and selecting strategies.
@@ 4.4 Effect of LLMs
Finding 5. There is no single LLM that is uniformly dominant across all strategies. Figure 4 shows the efficacy of proposed strategies for open-source LLMs, with detailed results in Appendix C. We can see that the F1 scores of the matching, comparing, and selecting strategies for different LLMs mostly fall between 50%~70%, 60%~80%, and 70%~80%, respectively. In general, similar to GPT-3.5 Turbo, the comparing strategy is better than the matching strategy, while the selecting strategy is further better than the comparing strategy. The consistent performance between strategies confirms the effectiveness of incorporating record interactions in these ways. Concretely, some chat LLMs, such as Llama3-8B and Mistral-7B, produce numerous false positives and thus perform poorly with the matching strategy. Nevertheless, they achieve significant improvement and satisfactory performance by comparing and selecting strategies, respectively. Moreover, although Flan-T5-XXL and Flan-UL2 lag behind GPT-4o by about 5% F1 in the selecting strategy, we find that they perform quite well in the matching and comparing strategies. These task-tuned LLMs follow instructions better than chat-tuned LLMs, and can output only the requested labels instead of long-form responses, making it convenient to utilize label generation probabilities. In conclusion, there is a noticeable variance in the capabilities of different LLMs for a single strategy, and the efficacy of different strategies for a single LLM can also be significantly distinct.
@@
Finding 6. Matching strategy is better for ranking and filtering than comparing strategy. The superiority of Flan-T5 in the matching and comparing strategies leads us to explore the possibility of using it to rank and filter potential matches for the selecting strategy. As shown in Figure 5, the matching strategy outperforms the comparing strategy under different model parameter sizes, even though the latter performs $\mathcal{O}(n^{2})$ comparisons. The difference is small on Flan-T5-XL (3B) and Flan-T5-XXL (11B), but significant on smaller models. This may be due to the fact that these models are trained on many pairwise tasks, such as natural language inference and question answering, but few triplewise tasks. Therefore, in terms of effectiveness and efficiency, the matching strategy is more suitable for ranking and filtering potential matches.
Figure 6: Average F1, Precision, and Recall w.r.t. number of candidate retained ($k$) for further selection.
@@ 4.5 Ablation Study
We perform ablation studies on the number of candidate records for further identification. As shown in Figure 6, the performance of GPT-3.5 Turbo is highly variable, while that of GPT-4o Mini is relatively stable. These results suggest that as LLMs evolve, ComEM may become more robust to the number of potential matches retained.
@@ 5 Conclusion
In this paper, we investigate three strategies of LLMs for entity matching to bridge the gap between local matching and global consistency of ER. Our research shows that incorporating record interactions is essential for LLM-based entity matching. By examining the effect of broad LLMs on these strategies, we further design a ComEM framework that integrates the advantages of multiple strategies and LLMs. The effectiveness and cost-efficiency of ComEM highlight the importance of task decomposition and LLM composition, opening up new avenues for entity matching using LLMs.
@@
This study aims to investigate different strategies for LLM-based entity matching. We conducted thorough experiments with two commercial LLMs and eight open-source LLMs to provide a broad base for our analysis. The selection of LLMs is based on considerations of popularity, availability, and cost. Future research could explore whether similar findings hold as LLMs evolve and how performance changes relative to our results.
Since LLMs were trained on massive amount of web data, they are likely to have seen similar and same records, or even some matching results, even though the labels of the matches are stored separately. Nevertheless, the performance of these strategies is relatively consistent across 10 LLMs and varies greatly for the same LLM when using different strategies, highlighting that data exposure is not the determining factor in their effectiveness. In the future, it will be valuable to evaluate LLM-based entity matching on new or non-public data.
The investigation of different strategies was conducted using basic zero- or few-shot prompting, a simple and effective paradigm for applying LLMs. We could not ignore the role of potential advanced prompt engineering methods in improving the accuracy and robustness of LLMs. In addition, fine-tuning LLMs for better execution of different strategies is also a worthwhile direction.
@@
Acknowledgments
We sincerely thank the reviewers for their insightful comments and valuable suggestions. This work was supported by the Natural Science Foundation of China (No. 62122077, 62106251) and Beijing Natural Science Foundation (L243006).
Appendix A Global Consistency of ER
In this paper, we refer to the interdependence of matching decisions in entity resolution as global consistency. This means that whether two records match is not an isolated decision, but is influenced by the matching results of other record pairs. This includes properties such as: 1) Reflexive: A record always matches itself; 2) Symmetric: If record A matches record B, then B also matches A; 3) Transitive: If A matches B, and B matches C, then A should match C; 4) Mutually exclusive: In some cases such as clean-clean ER, if A matches B, it cannot match C. Global consistency motivates the incorporation of more record interactions for LLM-based entity matching, rather than just considering two records independently.
Appendix B Strategy Prompts
The prompts for different strategies of LLM-based entity matching used in this paper are presented in Table 5. To ensure fairness, the same prompts were used for all experimental LLMs. These prompts were constructed through a manual process of prompt engineering, which involved the testing and comparing of multiple variations to determine the most effective ones. In addition to the task description, we included specific response instructions such as “Answer "Yes" if they do and "No" if they do not” to guide the responses of LLMs. For in-context learning, prompts and labels were repeatedly inputted for each example, followed by the records to be matched. We post-processed the LLM responses to obtain the final predicted labels.
Appendix C Detailed Results of Open-Source LLMs under Different Strategies
We provide the detailed F1 scores of open-source LLMs under different strategies and ComEM in Table 6. Among the eight LLMs evaluated in our experiment, six achieve the best performance through the selecting strategy, and two achieve better performance through the comparing strategy. In summary, our proposed strategies are universally applicable across different LLMs for entity matching. We have observed that it is difficult to limit the output of many chat-tuned LLMs simply by prompts, which may affect their actual performance. Therefore, how to calibrate the label probabilities from the long-form responses of LLMs is also important for performance improvement.